OpenStudy (goldphenoix):

Of my books, 85% are new and the rest are used. Some are biographies, 70% of which are new. What is the ratio of the fraction of new books that are biographies to the fraction of used books that are biographies? A. 7:17 B: 14:17 C: 17:14 D: 17:7

OpenStudy (anonymous):

wait let me reanalyze lol

OpenStudy (anonymous):

please tell me you've discussed "trees" or something that looks like it

OpenStudy (goldphenoix):

I don't think so. The only tree that I can think of is the one used in probability.

OpenStudy (anonymous):

how about the set theory thingy in statistics? the ones with the \(A\cap B\) familiar?

OpenStudy (anonymous):

because those things are the only ways i see to solve this

OpenStudy (goldphenoix):

I'm starting 8th grade this September. I'm not sure what you are talking about.

OpenStudy (anonymous):

darn. i can't narrow down my thought process

OpenStudy (anonymous):

maybe @terenzreignz will think of something before I do

OpenStudy (goldphenoix):

Help me TJ! D:

terenzreignz (terenzreignz):

Still thinking...

OpenStudy (goldphenoix):

I got this from the 2012-2013 Annual 7th Grade contest. I took it, and I did pretty bad... >_> I'm just checking how I can solve these problems, so that I can solve it later on.

OpenStudy (anonymous):

hmm... books... 85% new...15% used bios...70% new...30% used new books bio/ used books bio hmm... let me post this...it helps my thought process

OpenStudy (goldphenoix):

Alright. I wish I have your brain. I try everything in my head. Well, I kind of have to in this contest because there's no room! T~T

terenzreignz (terenzreignz):

Okay... I have an idea.

terenzreignz (terenzreignz):

Fraction of new books that are biographies = number of new biographies OVER number of new books.

OpenStudy (goldphenoix):

Is it (70/85)?

OpenStudy (goldphenoix):

^I doubt that's the answer though.

terenzreignz (terenzreignz):

Strictly speaking, it's \[\Large \frac{0.7 \times <number \ of \ biographies>}{0.85}\]

OpenStudy (anonymous):

if simplfied it matches one of the choices though

terenzreignz (terenzreignz):

And no, that's not the answer yet.

OpenStudy (goldphenoix):

I meant that's not number of new biographies over number of new books.

terenzreignz (terenzreignz):

But that <number of biographies> will be irrelevant once you consider the other fraction: Fraction of used books that are biographies = number of used biographies OVER number of used books

terenzreignz (terenzreignz):

Sorry... it would be \[\Large \frac{0.3\times <number \ of \ biographies>}{0.15}\]

OpenStudy (anonymous):

that doesn't match the choice though...i think it gives 7/16

terenzreignz (terenzreignz):

Patience, people :D

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

it was a typographical error...it does match

terenzreignz (terenzreignz):

ahahaha @xlegendx ;)

OpenStudy (goldphenoix):

My head hurt. T~T I do not understand what you mean by number of biographies.

terenzreignz (terenzreignz):

I mean, it wasn't specified, so the actual fraction would be (for the second one) 30% OF THE BIOGRAPHIES divided by the total number of used books :)

terenzreignz (terenzreignz):

Don't worry, it's just an unknown number that will prove irrelevant once we get to the next step... anytime you're ready :D

OpenStudy (goldphenoix):

I understand.

terenzreignz (terenzreignz):

Okay, great :) We now have the two fractions, and we need the ratio between them. Let's let me have an easier time and represent the <number of biographies> simply as b, shall we? \[\Large \frac{0.7\times b}{0.85}:\frac{0.3\times b}{0.15}\]

OpenStudy (goldphenoix):

Okay. :)

terenzreignz (terenzreignz):

So, that means b is a common factor of the left part and the right part of the ratio, it means it may be cancelled out :D

terenzreignz (terenzreignz):

Got that part? That's about the last tricky part of this question.

OpenStudy (goldphenoix):

Hmm. I believe so.

terenzreignz (terenzreignz):

To illustrate, the ratio \[2\times \color{red}3 : 3\times \color{red}3\] is the same as the ratio \[2 : 3\] ie the common factor of 3 may be cancelled out. It's like reducing to lowest terms, yeah? :D

OpenStudy (goldphenoix):

Yeah!

terenzreignz (terenzreignz):

Okay, let's get this done... cancelling out the common factor b, we get the "not so" simple ratio \[\Large \frac{0.70}{0.85}: \color{red}{\frac{0.30}{0.15}}\]

terenzreignz (terenzreignz):

We could multiply both sides of the ratio by \[\Large \frac{100}{100}=1\]To make things simpler...

OpenStudy (goldphenoix):

I see.

terenzreignz (terenzreignz):

Giving us... \[\Large \frac{70}{85}:\frac{30}{15}\] Carry on from here?

OpenStudy (goldphenoix):

Hmm. 70 * 15 =1050 85 * 30 = 2550 1050:2550 = 7:17

terenzreignz (terenzreignz):

Awww yeaaahh :D

OpenStudy (goldphenoix):

:D

OpenStudy (goldphenoix):

Is this suppose to be a 7th grade level?!

terenzreignz (terenzreignz):

I'm pretty sure I did something like this back in 6th :3

OpenStudy (goldphenoix):

Wowow! This seem so difficult. I've done this before.